# Beyond Likes: How Normative Feedback Complements Engagement Signals on Social Media

### v2

@@ Abstract.
Many online platforms incorporate engagement signals—such as likes and upvotes—into their content ranking systems and interface design. These signals are designed to boost user engagement. However, they can unintentionally elevate content that is less inclusive and may not support normatively desirable behavior. This issue becomes especially concerning when toxic content correlates strongly with popularity indicators such as likes and upvotes. In this study, we propose structured prosocial feedback as a complementary signal to likes and upvotes—one that highlights content quality based on normative criteria to help address the limitations of conventional engagement signals. We begin by designing and implementing a machine learning feedback system powered by a large language model (LLM), which evaluates user comments based on principles of positive psychology, such as individual well-being, constructive social media use, and character strengths. We then conduct a pre-registered user study to examine how existing peer-based and the new expert-based feedback interact to shape users’ selection of comments in a social media setting. Results show that peer feedback increases conformity to popularity cues, while expert feedback shifts preferences toward normatively higher-quality content. Moreover, incorporating expert feedback alongside peer evaluations improves alignment with expert assessments and contributes to a less toxic community environment. This illustrates the added value of normative cues—such as expert scores generated by LLMs using psychological rubrics—and underscores the potential benefits of incorporating such signals into platform feedback systems to foster healthier online environments.
@@
ccs: Human-centered computing Collaborative and social computing
@@ 1. Introduction
To assess the behavioral impact of this approach, we conduct a controlled user experiment in which participants complete content selection tasks while receiving real-time prosocial feedback on candidate posts. Our study is designed to test how this new feedback signal interacts with existing traditional popularity cues, and whether it meaningfully shifts users’ preferences or attention. In particular, we examine whether users are willing to prioritize value-aligned content even when it receives fewer likes or retweets.
@@ 2.4. Reimagining Feedback: From AI Training to Human Nudging
This formative potential of feedback is also recognized in community governance practice. In a recent study, Lambert et al. (Lambert et al., 2024) interviewed Reddit moderators and found that many deliberately provide positive feedback to encourage desired behavior to encourage desirable behaviors and foster a more constructive discourse environment. Yet, despite its perceived efficacy, moderators noted a lack of explicit infrastructural support for such feedback, revealing an opportunity for design interventions that systematize and amplify positive reinforcement in online communities.
Our work builds on this premise but reimagines the design space. Inspired by existing systems like Rotten Tomatoes which provides both peer (user ratings) and expert (Critic scores) reviews of movies, we use AI-based critics to encourage constructive posts instead of sanctioning less desirable ones.
@@ 4. User Study: Exploring Responses to Different Feedback Signals
To investigate the practical effects of our newly introduced feedback, we conducted a controlled user study simulating a comment-posting experience on a social media platform. We created an interactive online experiment hosted on Qualtrics, which mirrors a social media environment. Participants were asked to choose which comment they would prefer to post in response to a series of discussion questions (Fig 1). We opted for a selection task rather than asking participants to write their own responses in order to ensure consistent effort across all possible choices and to enable controlled comparisons. This design allows us to isolate the effect of feedback on participants’ judgments without introducing variance due to writing ability or motivation. Crucially, the comments shown were drawn from the Reddit dataset containing real user responses and were pre-selected to reflect varying feedback signals, enabling us to observe user preferences under different feedback conditions. The study was preregistered at https://aspredicted.org/msrh-xhqd.pdf.
Figure 1. User interface of our comment-posting simulation task. The left panel introduces the discussion topic question to participants. The right panel displays a set of pre-written comments, from which participants are asked to choose the one they would most like to post. Each comment is accompanied by a Peer Score and an Expert Score in the presented condition, simulating feedback signals to investigate their influence on user choices.
[…]
TBD
@@ 4.1. Constructing Feedback-Driven Comment Sets
    To simulate the dual feedback ecology of real-world platforms: Since social media platforms have long relied on popularity-based feedback, such as likes, upvotes, or shares, to gauge user engagement and influence content visibility, our newly introduced AI-generated, value-aligned feedback should be designed to coexist with these existing popularity-based signals. In this context, it is crucial to understand how users navigate the interplay between popularity-based feedback and our prosocial feedback.
@@ 4.2. Data Preparation
Figure 2. Main workflow diagram of the user study process.
[…]
TBD
@@ 4.3. Experimental Procedure
    Expert-Only Feedback: Only expert score is displayed.
@@
Finally, each participant selected a response to each question, from one of four choices. We use the Peer Score and Expert Score of this response in our analysis that follows.
We recruited 546 U.S. participants via Prolific, of whom 496 passed the attention check and were included in our analysis. As shown in Fig 2, participants first went through an onboarding phase, in which they were introduced to the overall structure of the study, including the decision task and the possible feedback mechanisms (peer score, expert score, or both). The purpose of this phase was to ensure participants had a clear understanding of the task and scoring systems before proceeding.
@@
Participants were instructed to select the comment they most preferred to post, with no right or wrong answers. After completing the four main decision tasks, participants answered a short questionnaire regarding their demographic information.
@@ 5.2.2. Permutation Test Validation
Peer-Only Feedback $0$ 0.000 0.000
Expert-Only Feedback 0.000 $0$ 0.000
Dual Feedback 0.463 0.000 $0$
@@
Peer-Only Feedback $0$ 0.000 0.000
Expert-Only Feedback 0.000 $0$ 0.000
Dual Feedback 0.495 0.000 $0$
@@ 6.4. When More Is Not Always Enough: The Challenges of Mixed Feedback
6.5. General Implications and Future Work
This work contributes to a growing conversation about how platforms can move beyond engagement-centric designs toward value-aligned interfaces, ones that encourage thoughtful participation, elevate high-quality contributions, and balance social and normative cues. Our results suggest that platforms do not necessarily need to abandon familiar peer-driven mechanisms to introduce expert guidance. Instead, they can incrementally enrich the feedback environment by integrating expert signals in ways that complement, rather than disrupt, existing social signals.
Looking ahead, several avenues for future research emerge. First, longitudinal studies could explore whether the effects we observed persist over time or change as users become more familiar with mixed feedback formats. Second, there’s a need to study how different framing techniques—such as narrative explanations, confidence indicators, or value-oriented prompts—can scaffold better understanding and use of evaluative signals. Finally, given that our study was conducted in a controlled lab setting, future work could involve A/B testing on actual platforms. Such deployment-based evaluations would help determine whether the observed effects generalize at scale and under real-world constraints.
By surfacing both the opportunities and tensions of mixed feedback design, our study provides a foundation for building platforms that support more informed, reflective, and socially constructive online participation.
@@ 7. Conclusion
This study investigated how different forms of evaluative feedback—peer-based, expert-based, and a combination of both—shape user selection behavior in online content environments. Our findings reveal that even lightweight feedback signals can meaningfully influence user choices: peer feedback promotes conformity to popularity, expert feedback nudges toward normatively strong content, and layering expert cues atop peer signals can augment their impact.
These insights have direct implications for the design of digital platforms seeking to promote higher-quality engagement. Rather than relying solely on engagement metrics, platforms can embed expert-aligned feedback in ways that guide users toward more thoughtful or prosocial contributions. Our work also suggests that thoughtful layering, contextual framing, and explanatory design elements are key to unlocking the full potential of mixed feedback systems.
@@
By clarifying how different types of feedback shape user decisions, this research lays a foundation for developing more value-aligned, cognitively supportive social systems. Future work may build on these findings by exploring adaptive, personalized feedback strategies, conducting longitudinal field studies, and extending these insights to diverse content domains.

### v3

@@ Abstract.
Many online platforms incorporate engagement signals—such as likes—into their interface design to boost engagement. However, these signals can unintentionally elevate content that may not support normatively desirable behavior, especially when toxic content correlates strongly with popularity indicators. In this study, we propose structured prosocial feedback as a complementary signal, which highlights content quality based on normative criteria. We design and implement an LLM-based feedback system, which evaluates user comments based on principles from positive psychology, such as individual well-being. A pre-registered user study then examines how existing peer-based (popularity) and the new expert-based feedback interact to shape users’ reposting behavior in a social media setting. Results show that peer feedback increases conformity to popularity cues, while expert feedback shifts choices toward normatively higher-quality content. This illustrates the added value of normative cues and underscores the potential benefits of incorporating such signals into platform feedback systems to foster healthier online environments.
@@
$c$s: Human-centered computing Collaborative and social computing
@@ 1. Introduction
To assess the behavioral impact of this approach, we conduct a controlled user experiment in which participants complete reposting tasks while receiving real-time prosocial feedback on candidate posts. Our study is designed to test how this new feedback signal interacts with existing traditional popularity cues, and whether it meaningfully shifts users’ preferences or attention. In particular, we examine whether users are willing to prioritize value-aligned content even when it receives fewer likes or retweets.
@@ 2.4. Reimagining Feedback: From AI Training to Human Nudging
This formative potential of feedback is also recognized in community governance practice. In a recent study, Lambert et al. (Lambert et al., 2024) interviewed Reddit moderators and found that many deliberately provide positive feedback to encourage desired behavior and foster a more constructive discourse environment. Yet, despite its perceived efficacy, moderators noted a lack of explicit infrastructural support for such feedback, revealing an opportunity for design interventions that systematize and amplify positive reinforcement in online communities.
Our work builds on this premise but reimagines the design space. Inspired by existing systems like Rotten Tomatoes, which provides both peer (user ratings) and expert (Critic scores) reviews of movies, we use AI-based critics to encourage constructive posts instead of sanctioning less desirable ones.
@@ 4. User Study: Exploring Responses to Different Feedback Signals
To investigate the practical effects of our newly introduced feedback, we conducted a controlled user study simulating a comment-reposting experience on a social media platform. We created an interactive online experiment hosted on Qualtrics, which mirrors a social media environment. Participants were asked to choose which comment they would prefer to repost in response to a series of discussion questions (Fig 1). We opted for a reposting task rather than asking participants to write their own responses in order to ensure consistent effort across all possible choices and to enable controlled comparisons. This design allows us to isolate the effect of feedback on participants’ judgments without introducing variance due to writing ability or motivation. Crucially, the comments shown were drawn from the Reddit dataset containing real user responses and were pre-selected to reflect varying feedback signals, enabling us to observe user preferences under different feedback conditions. The study was preregistered at https://aspredicted.org/msrh-xhqd.pdf.
Figure 1. User interface of our comment-reposting simulation task. The left panel introduces the discussion topic question to participants. The right panel displays a set of pre-written comments, from which participants are asked to choose the one they would most like to repost. Each comment is accompanied by a Peer Score and an Expert Score in the presented condition, simulating feedback signals to investigate their influence on user choices.
@@ 4.1. Constructing Feedback-Driven Comment Sets
    To simulate the dual feedback ecology of real-world platforms: Since social media platforms have long relied on popularity-based feedback, such as likes, upvotes, or shares, to gauge user engagement and influence content visibility, our newly introduced AI-generated, value-aligned feedback should be designed to coexist with these existing popularity-based signals. In this context, it is crucial to understand how users navigate the interplay between popularity-based (peer) feedback and our prosocial (expert) feedback.
@@ 4.3. Experimental Procedure
    Expert-Only Feedback: Only expert feedback is displayed.
@@
We recruited 546 U.S. participants via Prolific, of whom 496 passed the attention check and were included in our analysis. As shown in Fig 2, participants first went through an onboarding phase, in which they were introduced to the overall structure of the study, including the decision task and the possible feedback mechanisms (no feedback, peer feedback, expert feedback, or both). In particular, participants were given detailed explanations of how the two types of feedback were generated: the Peer Score estimated the content’s potential popularity among peers, while the Expert Score assessed the substantive quality of the content. The purpose of this phase was to ensure participants had a clear understanding of the task and scoring systems before proceeding.
@@
Participants were instructed to select the comment they most preferred to post, with no right or wrong answers. In this setup, participants did not generate new responses; instead, they effectively ‘reposted’ one of the existing comments as their response to the topic question. We use the Peer Score and Expert Score of this response in our analysis that follows. After completing the four main decision tasks, participants answered a short questionnaire regarding their demographic information.
Figure 2. Main workflow diagram of the user study process.
@@ 5.2.2. Permutation Test Validation
Peer-Only Feedback 0.000 - 0.000 0.000
Expert-Only Feedback 0.000 0.000 - 0.000
Dual Feedback 0.463 0.000 0.000 -
@@
Peer-Only Feedback 0.000 - 0.000 0.000
Expert-Only Feedback 0.000 0.000 - 0.000
Dual Feedback 0.495 0.000 0.000 -
@@ 6.4. When More Is Not Always Enough: The Challenges of Mixed Feedback
6.5. General Implications
This work contributes to a growing conversation about how platforms can move beyond engagement-centric designs toward value-aligned interfaces, ones that encourage thoughtful participation, elevate high-quality contributions, and balance social and normative cues. Our results suggest that platforms do not necessarily need to abandon familiar peer-driven mechanisms to introduce expert guidance. Instead, they can incrementally enrich the feedback environment by integrating expert signals in ways that complement, rather than disrupt, existing social signals. By surfacing both the opportunities and tensions of mixed feedback design, our study provides a foundation for building platforms that support more informed, reflective, and socially constructive online participation.
6.6. Limitations and Future Work
Our work has several limitations that point toward promising avenues for future work. First, the study was conducted in a controlled lab setting rather than in real-world settings. While this design enabled systematic comparisons across conditions, it may not fully reflect the dynamics of posting behaviors on live social media platforms. Deployments in the field, such as A/B tests on actual platforms, would help determine whether the observed effects generalize at scale and under real-world constraints. Second, our study focused on short-term behavioral outcomes, leaving open whether effects persist over time or shape longer-term attitudes and trust. Longitudinal studies are needed to assess the durability of these influences. Finally, our design presented expert feedback as a simple numerical score without explanations. Future designs could explore richer presentation formats, such as narrative rationales, confidence indicators, or value-oriented prompts, to help users interpret and integrate signals more effectively.
@@ 7. Conclusion
This study investigated how different forms of evaluative feedback—peer-based, expert-based, and a combination of both—shape user reposting behavior in online content environments. Our findings reveal that even lightweight feedback signals can meaningfully influence user choices: peer feedback promotes conformity to popularity, expert feedback nudges toward normatively strong content, and layering expert cues atop peer signals can augment their impact.
These insights have direct implications for the design of digital platforms seeking to promote higher-quality engagement. Rather than relying solely on engagement metrics, platforms can embed expert-aligned feedback in ways that guide users toward more thoughtful or prosocial contributions. Our work also suggests that thoughtful layering, contextual framing, and explanatory design elements could play an important role in unlocking the full potential of mixed feedback systems.
@@
By clarifying how different types of feedback shape user decisions, this research lays a foundation for developing more value-aligned, cognitively supportive social systems. Future work could build on these findings through longitudinal and real-world evaluations, as well as by exploring alternative definitions and designs of normative feedback.